Jason has a packet of tomato seeds left over from last year. He plants 36 of the seeds and only 8 sprout. what is the experimental probability that a tomato seed from this packet will sprout?
Answer by stanbon(75887)

Jason has a packet of tomato seeds left over from last year. He plants 36 of the seeds and only 8 sprout. what is the experimental probability that a tomato seed from this packet will sprout?
----------
P(sprout) = 8/36 = 2/9
-----
Cheers,
Stan H.